# Tuning without Peeking: Provable Privacy and Generalization Bounds for LLM Post-Training

### v1

@@ D.2 Possible Extensions
Code and Package. As said, all experiments in this work have been run on the Nevergrad open source platform [46], which encompasses many BBO algorithms, along with specialized wizards that select among them based on problem-specific properties. Nevergrad also includes algorithms that are black-box optimizers without being comparison-based, such as [130, 131]. Since these operate with a finite number of bits per loss function, they might also fall under the theoretical framework of Sec. 4, though this is left for future work.
@@
An important assumption (5) is that the only thing that depends on the dataset $D$ is $choice_{1},\dots,choice_{b}$, and each $choice_{i}$ has at most $k_{i}$ possible values ($k_{i}:=alg.numCalls()$ is called the branching factor [115]). This is unusual in deep learning (DL), as in DL $compare$ is replaced by some function involving gradients (even Reinforce [132] or GRPO [45] return a gradient update), and $choice_{i}$ is replaced by a gradient, or, for reinforcement learning such as GRPO [45], the outputs and their rewards.
@@
Because our results are mainly useful when $k_{i}$ is small, this work focused on comparison-based algorithms. However, all results of Sec. 4 are true for large values of $k_{i}$. Let us check $k_{i}$ in the case of some classical methods for post-training. In BBoxER, the information flow from the data to the model is limited to roughly one bit per iteration, leading to at most $2^{b}$ possible outputs for a budget of $b$ steps. This limited capacity directly reflects the compression properties showcased in Equation 3 and Figure 2. In contrast, supervised fine-tuning updates the model using one gradient per mini-batch. For an 8B model with 16-bit precision, each mini-batch carries about 256B bits: a few mini-batches are sufficient to encode the entire content of Wikipedia. Reinforcement learning methods, such as GRPO, typically rely on model outputs and reward signals that would allow them to encode the full dataset within a single training epoch.
@@ Appendix E Overfitting Bounds and their Generalizations
In this Section we provide the theoretical generalization bounds for the deterministic (Sec. E.1) and randomized cases (Sec. E.2), followed by (Sec. E.3) strengthened bounds for the special cases of $(1,\lambda)$ and $(1+\lambda)$ BBO algorithms as described in App. D. We proceed by making our bounds explicit for various families of algorithms, based on Fournier and Teytaud [115]; in Sec. E.4 we obtain strengthened bounds for Bet-and-Run combinations like 3-OneFifth, MultiDisc, or 2-DE/D-CMA  (deployed in some of our experiments in Sec. 5), Sec. E.5 lays out bounds for population-based evolution strategies (i.e., $(\mu+\lambda)$ and $(\mu,\lambda)$ with $\mu>1$) like CMA-ES and D-CMA, and Sec. E.6 deals with differential evolution and particle swarm optimization. The key to our results is the branching factor $k_{i}$, i.e., the number of distinct behaviors of the algorithm that can arise in an iteration. Finally, Sec. E.7 derives explicit bounds as a function of the dataset size.
@@ E.1 Bounds for the Deterministic Algorithm: Theorem 1
For Equation A.4: apply the Bonferroni bound (Equation A.3) to the set of events
@@
For Equation A.5, apply the Bonferroni bound to the set of events
@@ E.2 Generalization Bounds for Randomized BBoxER
Thm 1 applies to deterministic algorithms, where the only source of randomness is the sampling of the dataset $D$, determined by the seed $\omega_{D}$. In this case, the deviation risk $\delta_{1,\epsilon}$ is computed over that randomness alone. When the optimization algorithm itself is randomized (i.e., $\omega$ is a random seed), the deviation risk $\delta_{\epsilon,\omega}$ also depends on $\omega$. Averaging over $\omega$ yields the total deviation risk:
@@
Since the bound of Thm 1 holds uniformly for any fixed $\omega$, it extends naturally to the randomized setting, in which $\widehat{x}$ depends on both $\omega$ and $\omega_{D}$:
@@
In contrast to random search, where all candidate parametrizations are chosen independently of their evaluation, comparison-based algorithms build candidates iteratively based on performance. If $\lambda$ candidates are evaluated at each step and the best is retained, then the number of internal states after one iteration is $\lambda$; after two iterations, $\lambda^{2}$; and after $n=b/\lambda$ iterations, at most $\lambda^{n}$. The total risk is, therefore, bounded by $\lambda^{n}\cdot\delta_{1,\epsilon}$, where $n$ is the number of iterations with $k_{i}>1$ (with $k_{i}=\lambda$ in such iterations); see [115].
In the elitist variant $(1+\lambda)$, the best model is selected among the current best and the $\lambda$ new candidates, so $k_{i}\in\{1,\lambda+1\}$. This leads to a total number of internal states bounded by $(\lambda+1)^{(b-1)/\lambda}$. Simply apply Thm 3 in the case
@@ E.4 Generalization Bounds for Bet-And-Run Combinations
Consider the example in Equation A.7. Additionally, if we perform $k>1$ independent optimization runs, and if the budget $b$ is divided by $k$, then the Bonferroni correction leads to a risk
$\underbrace{k}_{\text{Bonferroni correction}}\cdot\underbrace{\lambda^{(b/(\lambda\cdot k))}}_{\text{Branching factor per run}}\cdot\underbrace{\delta_{1,\epsilon}}_{\text{Risk for a single model}}$ (A.8)
for the resulting parametrization. According to these bounds, increasing the number $k$ of independent runs for a given total budget decreases the risk of overfitting. More generally, consider $bet-and-run(a_{1},a_{2},\dots,a_{k})$, the bet-and-run of $k$ algorithms $a_{1}$,…,$a_{k}$. This algorithm, for a budget $b$, runs each $a_{i}$ with budget $b/k$:
@@
Let us consider that the $a_{i}$ are all the same algorithm, and let us check the impact of bet-and-run on overfitting. Equation A.9 implies a lower overfitting than with a single algorithm (corresponding to $N(\omega,a_{i},b)$) if $N(\omega,a_{i},b/k)$ increases more than linearly in $b/k$, which is the case for most optimization algorithms. For example, with $(1+1)$ evolution strategies, Sec. E.3 shows that $N(\omega,evolution\_strategy,b)$ is exponential in $b$.
Let us illustrate this in an example. Consider Equation 5 in the case $k_{i}\leq 2$, namely:
@@
In the case of CMA or D-CMA, the $(\mu,\lambda)$ strategy is more usual and used in our experiments; and typically, the population-size $\lambda$ is set as $\lambda=4+\lfloor 3\log(d)\rfloor$ in dimension $d$, and $\mu$ frequently scales as $\lambda/4$ or $\lambda/2$. Sometimes there is a complete ranking of the $\mu$ selected points so that different weights are used for the $\mu$ selected points depending on their rank: this adds a factor $\mu!$ to the number of possible branches: otherwise, we get bounds as in Tab. 3.
@@ E.7 Dependence on Dataset Size: Scaling and Numerical Application
Sec. C.2 presents classical large deviation inequalities, and the present paper shows how, in the case of comparison-based algorithms, we can derive non-vacuous bounds. Let us now see numerically how our overfitting bound behaves in practical cases.
@@
We have seen in Equation A.2 that, using Hoeffding’s bound, $\delta_{1,\epsilon}=2\exp(-2s\epsilon^{2})$ with $s=|D|$ the sample size, and $\delta\leq 2\times(2^{b})\exp(-s\epsilon^{2})$, leading to $\delta$ constant when $b$ scales linearly with $s$: the acceptable number of iterations (i.e., a fixed bound on $\delta$) is linear in the sample size $D$. Applying Hoeffding, and aiming at a risk $\delta$, we get that overfitting at level $\epsilon$ is impossible if
@@
If we prefer the Bennet&Bernstein inequality (based on an assumption on the standard deviation, see Sec. C.1 [77, Section 8.2]), we get
@@
This equation, for $s=8000$ examples (close to GSM8k), $\delta=\frac{1}{2}$ (we are considering the median case), $\epsilon=0.01$ (we want that precision) and $\sigma=0.06$ (the standard variation of the loss is 0.06, we assume that the initial model is already good), leads to the claim that a budget $b\leq 91$ cannot lead to overfitting. If we assume $\epsilon=0.04$ and $\sigma=0.3$, we get with Hoeffding a bound $34$ and with Bennet&Bernstein a bound $88$. With $\epsilon=0.06$, $\sigma=0.3$, Bennet&Bernstein leads to a budget $189$. The budget becomes $482$ if we use $\epsilon=0.1$, $\sigma=0.3$. Also, these bounds scale linearly with the number $s$ of feedbacks (dataset size or number of feedbacks in A/B testings). These numbers are low, but we see that bounds are not vacuous even in realistic scenarios.
@@ F.1 Privacy by Design
In contrast, the BBoxER approach is an instance of privacy by design [138]: It does not rely on gradients, nor does it require access to the raw content of user sessions. Instead, it only leverages a binary satisfaction signal–whether a user prefers one version of the model over another. This can be collected via simple A/B testing, and the algorithm proceeds using aggregated comparison outcomes. In particular, our method operates without knowledge of the specific queries or model responses, which never leave the user’s device.
@@
$P(A(D_{1})\in S)\leq\exp(\epsilon)\cdot P(A(D_{2})\in S)+\delta.$ (A.11)
We now examine how this applies in two distinct cases: the privacy of user queries and responses and the privacy of satisfaction feedback.
@@
Our algorithm never observes the actual content of user interactions. If $D_{1}$ and $D_{2}$ differ by one prompt or answer or any data unrelated to preference signals, then the outcome of Alg. 1 remains unchanged. In that case, the algorithm behaves identically on both datasets, satisfying differential privacy with $\epsilon=0$, $\delta=0$. This guarantee is stronger than what is typically offered in federated learning: since no gradients are shared or aggregated, the system is inherently robust to inference attacks such as those in [8, 9, 4].
Privacy of Satisfaction Signals.
We may also be interested in the privacy implications of modifying a single satisfaction response $r_{i}$ associated with a specific user. While this is less sensitive than access to the user’s actual queries or responses, it still reflects individual feedback and may warrant protection. In this setting, we consider the case where $D_{1}$ and $D_{2}$ differ by exactly one such response, altered during a run.
Applying Equation 6 with $k=1$, we get that $A(D_{1})$ and $A(D_{2})$ are the same with probability at least $1-3b/\sqrt{2n\pi}$ when we have $n$ data points and $b$ iterations. This corresponds to differential privacy with $\epsilon=0$ and $\delta=3b/\sqrt{2n\pi}$.
@@
In supervised fine-tuning, both inputs and outputs are required for training. Reinforcement learning additionally requires access to the model’s responses and reward signals. While federated learning aims to limit privacy leakage by aggregating gradients locally, it still exposes sensitive intermediate data—such as gradients—that have been shown to be vulnerable to attacks [9]. In contrast, our approach bypasses these vulnerabilities entirely by never interacting with such information in the first place.
@@ F.2 Poisoning
Poisoning attacks seek to influence the outcome of a learning algorithm by injecting malicious data. This threat has been widely studied in machine learning [140, 41, 40], and it is particularly relevant in interactive settings such as retrofitting. In our context, a poisoning attempt manipulates up to $k$ user feedbacks per iteration to bias the optimization process.
Our robustness Thm 2 quantifies the impact of such attacks: the probability that the final model is altered remains bounded by $\frac{(2k+1)b}{\sqrt{2n\pi}}$. This implies that an adversary must control $k\sim\sqrt{n}/b$ feedbacks per round to affect the result significantly. Such an attack becomes increasingly difficult as the number of users $n$ grows.
@@ G.1 Proof of Thm 2
Consider a frequency $f=\frac{1}{n}\sum_{i=1}^{n}r_{i}$ computed by considering $n$ samples $r_{1},\dots,r_{n}$ independent and identically distributed with $P(r_{i}=1)=f_{0}$ and $P(r_{i}=0)=1-f_{0}$ for some unknown $f_{0}$. Then
$P_{r_{1},\dots,r_{n}}\left(\exists\ f^{\prime}\mbox{ such that $|f^{\prime}-f|\leq k/n$ \text{ and } }(f-\frac{1}{2})\cdot(f^{\prime}-\frac{1}{2})\leq 0\right)\leq\frac{2k+1}{\sqrt{2n\pi}}.$ (A.12)
@@
First, using [77, Lemma A.3], we observe that
$\forall x\in\{0,\ldots,n\},P(B(n,\frac{1}{2})=x)\leq\frac{1}{\sqrt{2n\pi}}.$ (A.13)
Consider a probability $p$ that individual comparisons are in favor of a model $m_{1}$ against a model $m_{2}$. Assume that a modification of the training set has an impact $\epsilon$ on the frequency $f$ with which $m_{1}$ was preferred against model $m_{2}$: the new frequency $f^{\prime}$ is in $[f-\epsilon,f+\epsilon]$. $(f-\nicefrac{{1}}{{2}})\times(f^{\prime}-\nicefrac{{1}}{{2}})\leq 0$ is possible only if $f\in[\nicefrac{{1}}{{2}}-\epsilon,\nicefrac{{1}}{{2}}+\epsilon]$. And since $f\sim\frac{1}{n}B(n,p)$, $P(f\in[\nicefrac{{1}}{{2}}-\epsilon,\nicefrac{{1}}{{2}}+\epsilon])$ is maximum if $p=\nicefrac{{1}}{{2}}$. Then,
$P\left(f\in[\nicefrac{{1}}{{2}}-\epsilon,\nicefrac{{1}}{{2}}+\epsilon]\right)$ $\leq P\left(\frac{1}{n}B(n,\nicefrac{{1}}{{2}})\in[\nicefrac{{1}}{{2}}-\epsilon,\nicefrac{{1}}{{2}}+\epsilon]\right)$
$\leq P\left(B(n,\nicefrac{{1}}{{2}})\in[n(\nicefrac{{1}}{{2}}-\epsilon),n(\nicefrac{{1}}{{2}}+\epsilon)]\right)$
@@
If someone modifies the data by modifying the answers corresponding to $k$ users, the frequency moves to $f^{\prime}$ instead of $f$ with $|f-f^{\prime}|\leq\epsilon=k/n$ where $n$ is the number of independent users. This shows Equation A.12. ∎
To show Thm 2, we apply the Bonferroni correction when $k$ data values are modified at each iteration and we have $k$ iterations. Then, the probability of a difference between $output_{1}$ and $output_{2}$ is upper bounded by $b\times\delta$. This shows Equation 6.
G.2 Theorem on Data Extraction and Proof of Corollary 2
Corollary 2 is a direct consequence of the following Theorem.

### v2

@@ D.2 Possible Extensions
Code and Package. As said, all experiments in this work have been run on the Nevergrad open source platform [32], which encompasses many BBO algorithms, along with specialized wizards that select among them based on problem-specific properties. Nevergrad also includes algorithms that are black-box optimizers without being comparison-based, such as [125, 126]. Since these operate with a finite number of bits per loss function, they might also fall under the theoretical framework of Sec.˜4, though this is left for future work.
@@
An important assumption (˜5) is that the only thing that depends on the dataset $D$ is $choice_{1},\dots,choice_{b}$, and each $choice_{i}$ has at most $k_{i}$ possible values ($k_{i}:=alg.numCalls()$ is called the branching factor [108]). This is unusual in deep learning (DL), as in DL comparecompare is replaced by some function involving gradients (even Reinforce [127] or GRPO [31] return a gradient update), and $choice_{i}$ is replaced by a gradient, or, for reinforcement learning such as GRPO [31], the outputs and their rewards.
@@
Because our results are mainly useful when $k_{i}$ is small, this work focused on comparison-based algorithms. However, all results of Sec.˜4 are true for large values of $k_{i}$. Let us check $k_{i}$ in the case of some classical methods for post-training. In BBoxER, the information flow from the data to the model is limited to roughly one bit per iteration, leading to at most $2^{b}$ possible outputs for a budget of $b$ steps. This limited capacity directly reflects the compression properties showcased in Equation˜3 and Figure˜2. In contrast, supervised fine-tuning updates the model using one gradient per mini-batch. For an 8B model with 16-bit precision, each mini-batch carries about 256B bits: a few mini-batches are sufficient to encode the entire content of Wikipedia. Reinforcement learning methods, such as GRPO, typically rely on model outputs and reward signals that would allow them to encode the full dataset within a single training epoch.
@@ Appendix E Overfitting Bounds and their Generalizations
In this Section we provide the theoretical generalization bounds for the deterministic (Sec.˜E.1) and randomized cases (Sec.˜E.2), followed by (Sec.˜E.3) strengthened bounds for the special cases of $(1,\lambda)$ and $(1+\lambda)$ BBO algorithms as described in App.˜D. We proceed by making our bounds explicit for various families of algorithms, based on Fournier and Teytaud [108]; in Sec.˜E.4 we obtain strengthened bounds for Bet-and-Run combinations like 3-OneFifth, MultiDisc, or 2-DE/D-CMA (deployed in some of our experiments in Sec.˜5), Sec.˜E.5 lays out bounds for population-based evolution strategies (i.e., $(\mu+\lambda)$ and $(\mu,\lambda)$ with $\mu>1$) like CMA-ES and D-CMA, and Sec.˜E.6 deals with differential evolution and particle swarm optimization. The key to our results is the branching factor $k_{i}$, i.e., the number of distinct behaviors of the algorithm that can arise in an iteration. Finally, Sec.˜E.7 derives explicit bounds as a function of the dataset size.
@@ E.1 Bounds for the Deterministic Algorithm: Theorem 1
For Equation˜A.4: apply the Bonferroni bound (Equation˜A.3) to the set of events
@@
For Equation˜A.5, apply the Bonferroni bound to the set of events
@@ E.2 Generalization Bounds for Randomized BBoxER
Thm˜1 applies to deterministic algorithms, where the only source of randomness is the sampling of the dataset $D$, determined by the seed $\omega_{D}$. In this case, the deviation risk $\delta_{1,\epsilon}$ is computed over that randomness alone. When the optimization algorithm itself is randomized (i.e., $\omega$ is a random seed), the deviation risk $\delta_{\epsilon,\omega}$ also depends on $\omega$. Averaging over $\omega$ yields the total deviation risk:
@@
Since the bound of Thm˜1 holds uniformly for any fixed $\omega$, it extends naturally to the randomized setting, in which $\widehat{x}$ depends on both $\omega$ and $\omega_{D}$:
@@
In contrast to random search, where all candidate parameterizations are chosen independently of their evaluation, comparison-based algorithms build candidates iteratively based on performance. If $\lambda$ candidates are evaluated at each step and the best is retained, then the number of internal states after one iteration is $\lambda$; after two iterations, $\lambda^{2}$; and after $n=b/\lambda$ iterations, at most $\lambda^{n}$. The total risk is, therefore, bounded by $\lambda^{n}\cdot\delta_{1,\epsilon}$, where $n$ is the number of iterations with $k_{i}>1$ (with $k_{i}=\lambda$ in such iterations); see [108].
In the elitist variant $(1+\lambda)$, the best model is selected among the current best and the $\lambda$ new candidates, so $k_{i}\in\{1,\lambda+1\}$. This leads to a total number of internal states bounded by $(\lambda+1)^{(b-1)/\lambda}$. Simply apply Thm˜3 in the case
@@ E.4 Generalization Bounds for Bet-And-Run Combinations
Consider the example in Equation˜A.7. Additionally, if we perform $k>1$ independent optimization runs, and if the budget $b$ is divided by $k$, then the Bonferroni correction leads to a risk
[…]
for the resulting parametrization. According to these bounds, increasing the number $k$ of independent runs for a given total budget decreases the risk of overfitting. More generally, consider $bet-and-run(a_{1},a_{2},\dots,a_{k})$, the bet-and-run of $k$ algorithms $a_{1}$$a_{k}$. This algorithm, for a budget $b$, runs each $a_{i}$ with budget $b/k$:
@@
Let us consider that the $a_{i}$ are all the same algorithm, and let us check the impact of bet-and-run on overfitting. Equation˜A.9 implies a lower overfitting than with a single algorithm (corresponding to $N(\omega,a_{i},b)$) if $N(\omega,a_{i},b/k)$ increases more than linearly in $b/k$, which is the case for most optimization algorithms. For example, with $(1+1)$ evolution strategies, Sec.˜E.3 shows that $N(\omega,evolution\_strategy,b)$ is exponential in $b$.
Let us illustrate this in an example. Consider Equation˜5 in the case $k_{i}\leq 2$, namely:
@@
In the case of CMA or D-CMA, the $(\mu,\lambda)$ strategy is more usual and used in our experiments; and typically, the population-size $\lambda$ is set as $\lambda=4+\lfloor 3\log(d)\rfloor$ in dimension $d$, and $\mu$ frequently scales as $\lambda/4$ or $\lambda/2$. Sometimes there is a complete ranking of the $\mu$ selected points so that different weights are used for the $\mu$ selected points depending on their rank: this adds a factor $\mu!$ to the number of possible branches: otherwise, we get bounds as in Tab.˜4.
@@ E.7 Dependence on Dataset Size: Scaling and Numerical Application
Sec.˜C.2 presents classical large deviation inequalities, and the present paper shows how, in the case of comparison-based algorithms, we can derive non-vacuous bounds. Let us now see numerically how our overfitting bound behaves in practical cases.
@@
We have seen in Equation˜A.2 that, using Hoeffding’s bound, $\delta_{1,\epsilon}=2\exp(-2s\epsilon^{2})$ with $s=|D|$ the sample size, and $\delta\leq 2\times(2^{b})\exp(-s\epsilon^{2})$, leading to $\delta$ constant when $b$ scales linearly with $s$: the acceptable number of iterations (i.e., a fixed bound on $\delta$) is linear in the sample size $D$. Applying Hoeffding, and aiming at a risk $\delta$, we get that overfitting at level $\epsilon$ is impossible if
@@
If we prefer the Bennet&Bernstein inequality (based on an assumption on the standard deviation, see Sec.˜C.1 [102, Section 8.2]), we get
@@
This equation, for $s=8000$ examples (close to GSM8k), $\delta=\frac{1}{2}$ (we are considering the median case), $\epsilon=0.01$ (we want that precision) and $\sigma=0.06$ (the standard variation of the loss is 0.06, we assume that the initial model is already good), leads to the claim that a budget $b\leq 91$ cannot lead to overfitting. If we assume $\epsilon=0.04$ and $\sigma=0.3$, we get with Hoeffding a bound $34$ and with Bennet&Bernstein a bound $88$. With $\epsilon=0.06$, $\sigma=0.3$, Bennet&Bernstein leads to a budget $189$. The budget becomes $482$ if we use $\epsilon=0.1$, $\sigma=0.3$. Also, these bounds scale linearly with $s$ (the number of user preferences in A/B testings or dataset size). These numbers are low, but we see that our bounds are not vacuous even in realistic scenarios.
@@ F.1 Privacy by Design
In contrast, the BBoxER approach is an instance of privacy by design [133]: It does not rely on gradients, nor does it require access to the raw content of user sessions. Instead, it only leverages a binary satisfaction signal–whether a user prefers one version of the model over another. This can be collected via simple A/B testing, or from the individual samples of the dataset by comparing the models over the singleton {sample}. Then the algorithm proceeds using aggregated comparison outcomes. In particular, for A/B testing, our method operates without knowledge of the specific queries or model responses, that do not need to leave the user’s device, and in the general case, we only feed the algorithm the aggregate of model comparison without direct access to the underlying data.
@@
$P(A(D_{1})\in S)\leq\exp(\varepsilon)\cdot P(A(D_{2})\in S)+\delta.$ (A.11)
We now examine how this applies in two distinct cases: the privacy of user queries and responses and the general case of privacy.
@@
(like in A/B testing setting) Our algorithm never observes the actual content of user interactions. If $D_{1}$ and $D_{2}$ differ by one prompt or answer or any data unrelated to preference signals, then the outcome of Alg.˜1 remains unchanged. In that case, the algorithm behaves identically on both datasets, satisfying differential privacy with $\varepsilon=0$, $\delta=0$. This guarantee is stronger than what is typically offered in federated learning: since no gradients are shared or aggregated, the system is inherently robust to inference attacks such as those in [6, 7, 4].
Privacy of the entire sample.
(general case of applying BBoxER on a dataset $D$) We may also be interested in the privacy implications of modifying a single satisfaction response $r_{i}$ associated with a specific user, or in the case were we compute this preference signal from the data. While this is less sensitive than access to the user’s actual queries or responses, it still reflects user preference and may warrant protection. In this setting, we consider the case where $D_{1}$ and $D_{2}$ differ by exactly one such response, altered during a run.
Applying Equation˜7 with $m=1$, we get that $A(D_{1})$ and $A(D_{2})$ are the same with probability at least $1-3b/\sqrt{2s\pi}$ when we have $s$ data points and $b$ iterations. This corresponds to differential privacy with $\varepsilon=0$ and $\delta=3b/\sqrt{2s\pi}$.
@@
In supervised fine-tuning, both inputs and outputs are required for training. Reinforcement learning additionally requires access to the model’s responses and reward signals. While federated learning aims to limit privacy leakage by aggregating gradients locally, it still exposes sensitive intermediate data—such as gradients—that have been shown to be vulnerable to attacks [7]. In contrast, our approach bypasses these vulnerabilities entirely by never interacting with such information in the first place, in the case of A/B testing. Or, by limiting the impact of individual samples through an aggregate compressed signal from the data.
@@ F.2 Poisoning
Poisoning attacks seek to influence the outcome of a learning algorithm by injecting malicious data. This threat has been widely studied in machine learning [135, 28, 136], and it is particularly relevant in interactive settings such as retrofitting. In our context, a poisoning attempt manipulates up to $m$ user preferences per iteration to bias the optimization process.
Our robustness Thm˜2 quantifies the impact of such attacks: the probability that the final model is altered remains bounded by $\frac{(2m+1)b}{\sqrt{2s\pi}}$. This implies that an adversary must control $m\sim\sqrt{s}/b$ user preferences per round to affect the result significantly. Such an attack becomes increasingly difficult as the number of users (or dataset size) $s$ grows.
@@ G.1 Proof of Thm˜2
Consider a frequency $f=\frac{1}{s}\sum_{i=1}^{s}r_{i}$ computed by considering $s$ samples $r_{1},\dots,r_{s}$ independent and identically distributed with $P(r_{i}=1)=f_{0}$ and $P(r_{i}=0)=1-f_{0}$ for some unknown $f_{0}$. Then
$P_{r_{1},\dots,r_{s}}\left(\exists\ f^{\prime}\mbox{ such that $|f^{\prime}-f|\leq m/s$ \text{ and } }(f-\frac{1}{2})\cdot(f^{\prime}-\frac{1}{2})\leq 0\right)\leq\frac{2m+1}{\sqrt{2s\pi}}.$ (A.12)
@@
First, using [102, Lemma A.3], we observe that
$\forall x\in\{0,\ldots,s\},P(B(s,\frac{1}{2})=x)\leq\frac{1}{\sqrt{2s\pi}}.$ (A.13)
Consider a probability $p$ that individual comparisons are in favor of a model $model_{1}$ against a model $model_{2}$. Assume that a modification of the training set has an impact $\epsilon$ on the frequency $f$ with which $model_{1}$ was preferred against model $model_{2}$: the new frequency $f^{\prime}$ is in $f-\epsilon,f+\epsilon]$. $(f-\nicefrac{{1}}{{2}})\times(f^{\prime}-\nicefrac{{1}}{{2}})\leq 0$ is possible only if $f\in[\nicefrac{{1}}{{2}}-\epsilon,\nicefrac{{1}}{{2}}+\epsilon]$. And since $f\sim\frac{1}{s}B(s,p)$, $P(f\in[\nicefrac{{1}}{{2}}-\epsilon,\nicefrac{{1}}{{2}}+\epsilon])$ is maximum if $p=\nicefrac{{1}}{{2}}$. Then,
$P\left(f\in[\nicefrac{{1}}{{2}}-\epsilon,\nicefrac{{1}}{{2}}+\epsilon]\right)$ $\leq P\left(\frac{1}{s}B(s,\nicefrac{{1}}{{2}})\in[\nicefrac{{1}}{{2}}-\epsilon,\nicefrac{{1}}{{2}}+\epsilon]\right)$
$\leq P\left(B(s,\nicefrac{{1}}{{2}})\in[s(\nicefrac{{1}}{{2}}-\epsilon),s(\nicefrac{{1}}{{2}}+\epsilon)]\right)$
@@
If someone modifies the data by modifying the answers corresponding to $m$ users, the frequency moves to $f^{\prime}$ instead of $f$ with $|f-f^{\prime}|\leq\epsilon=m/s$ where $s$ is the number of independent users/samples. This shows Equation˜A.12. ∎
To show Thm˜2, we apply the Bonferroni correction when $m$ data values are modified at each iteration and we have $b$ iterations. Then, the probability of a difference between $output_{1}$ and $output_{2}$ is upper bounded by $b\times\delta$. This shows Equation˜7.
G.2 Theorem on Data Extraction and Proof of Corollary˜2
Corollary˜2 is a direct consequence of the following Theorem.